lim x tends to 90 deegree


sin(x cosx)/cos(x sinx)

Grade:12

3 Answers

Gurwinder Kaur
13 years ago

zero...

deepak kumar
13 years ago

how its zero

Gurwinder Kaur
12 years ago

bcoz,  sin(90cos90)/cos(90sin90) = sin(90*0)/cos(90*1)= sin0/cos90 = zero..